# Seamless Integration: Sampling Strategies in Federated Learning Systems

### v1

@@ II State of the Art
Figure 1: Illustration of the Federated Learning process [19]
@@ III Methodology
First, we examined how clustering approaches, which were primarily demonstrated on MNIST and minimal examples, can be transferred to real-world datasets and state-of-the-art neural networks, such as EfficientNet or DenseNet. Subsequently, we evaluated various clustering methods and selected the most effective one for further analysis. Initial testing was conducted on a somewhat simpler, proprietary dataset to establish a baseline understanding, which was then validated on a subset of the more complex ImageNet dataset. This approach allowed us to assess the generalizability and robustness of the selected clustering method across different types of data and neural network architectures.
Even relatively small neural networks typically contain millions of trainable parameters. When fine-tuning on custom datasets, it is common practice to adapt only the final layers to the specific data and the required number of output classes. Consequently, analyzing only the final layer’s activations is often sufficient for effective clustering, as it captures the most relevant features learned during fine-tuning. In this case, with a shape of (n, 1280) corresponding to n classes, this approach results in a small sample size with high dimensionality, which presents unique challenges for clustering.
@@ III-A Dataset
Our Dataset consists of cabins of miniature trucks, where the parts can be 3D printed, milled, or bought separately. See Figure 2 for two exemplary images from the quality inspection module. As the product is assembled to the wishes of the customer in lot size one, many combinations are possible. Previous publications proved that one advantage of FL is the higher generalizability [43], therefore allowing us to take many pictures away from the production line, where a higher variability of image date is easier to obtain and the risk of interfering with the production process is avoided. For testing purposes, we split the data set so that each client has only one color and one type of windshield, in addition to the class without windshield for every client. The distribution can be obtained from Table I. For some tests, not all clients take part in the federation from the beginning and simulate newcomers that join with new data.
@@ III-B Usability study on clustering
A pretrained EfficientNet B4 [44] was used as the frozen backbone, with one fully connected layer replaced to match the number of classes. The training was done with SGD as the optimizer with a learning rate of 0.001, 200 communication rounds and 5 local epochs. Using Pytorch, the preliminary examinations for clustering were carried out with our implementation to exclude any other influences (e.g. calculating the average manually based on client’s weight files).
To determine which clustering method is most appropriate, three methods that are well-suited for a large number of data points and clusters were selected: Agglomerative clustering, HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) [45] and BIRCH (Balanced Iterative Reducing and Clustering using Hierarchies) [46]. Figure 3 presents sample results from a larger set of tests that were conducted. The optimal parameters, such as the distance metric, for each method were determined through grid search on the weights from previous runs in which all clients participated. We observe that with agglomerative clustering, accuracy increases rapidly, particularly during the first 10 rounds, and then remains stable while continuing to improve gradually. All other methods exhibit significantly higher fluctuations and, except for BIRCH (n=5), even achieve lower accuracy than agglomerative clustering. It should be noted that BIRCH occasionally even clearly surpasses it. The high fluctuation in BIRCH’s performance could be due to its sensitivity to parameters and the high dimensionality combined with a small sample size. BIRCH is generally designed for larger datasets, and when applied to a small dataset with high dimensional features, it may struggle to form stable clusters, leading to the observed variability in test accuracy. HDBSCAN consistently performed worse in comparison.
Figure 4: Enter Caption
@@ III-C Evaluation of Client Selection Strategies
In order to compare different strategies, the clients as listed in Table II were selected for training and the data of the excluded clients were used for testing, whereby random images were selected to create a similar distribution of classes, e.g. to avoid over-representing class 0. The total number of clients was 22, which corresponds to ”all”. 16 clients correspond to the setup shown in Table II. In order to simulate the complete absence of a class, the last client with class 5 was omitted from the training in some tests, resulting in 15 clients. The configuration for each run was: EfficientNet with SGD as an optimizer, 200 communication rounds with 5 local epochs each. As a baseline, random sampling of approximately twenty and fifty percent (rounding up to the next integer number) of all participating clients was done. Between a sampling rate of 50% and 20% of all 22 clients, there is no noticeable difference, supporting the findings of [18] that saturation occurs at a certain point, and additional client participation no longer results in improvement.
@@
Figure 6 illustrates the test accuracy over 200 communication rounds for four client selection methods: Random Selection, Training Loss-based Selection, Cluster Selection with k=3, and Cluster Selection with k=5. The moving average is used for clearer visualization. Cluster Selection with k=5 consistently outperforms the other methods, showing the highest and most stable accuracy throughout. Initially, it demonstrates a rapid increase in accuracy, maintaining superior performance across all phases. Cluster Selection with k=3 and Random Selection yield similar results, with k=3 exhibiting more fluctuations but eventually aligning with Random Selection. The Training Loss method, despite early instability, eventually achieves comparable accuracy, though with more fluctuations. Overall, clustering with k=5 offers the greatest accuracy and stability, while non-randomized methods provide better and more controlled integration of new clients.
@@
Figure 5: Comparison of client participation frequencies across different selection methods: The chart displays the number of participations for each client using three selection strategies: Random Selection, Highest Training Loss and Cluster Selection with k=3 ”
Figure 6: The plot shows the moving average of test accuracy (with a window size of 5) across 200 communication rounds for three client selection methods: Random Selection (green), Training Loss-based Selection (blue), Cluster Selection with k=3 (dark red), and Cluster Selection with k=5 (red).
To ensure that cluster-based client selection is effective beyond our dataset, we conducted further experiments using a subset of ImageNet. This subset includes three superclasses: animals, tools, and vehicles, each containing multiple subclasses (for instance, various classes for cats, dogs, and birds under the animals superclass). Throughout all runs, the mapping of clients to datasets remained consistent, as did the randomization seed, such as when adding noise to the initially loaded ImageNet weights. The results, shown in Figure 7, demonstrate that clustering strategies with any amount of clusters outperform the random and training loss-based strategies, particularly during the first 25 communication rounds. This analysis was conducted using the same hyperparameters as before, highlighting the effectiveness of clustering approaches in improving model accuracy, especially in the early stages of training, but also proofing that ImageNet is more complex and needs more fine-tuning to achieve a higher accuracy. It is important to note that clustering with k=3 outperforms random selection with n=5. This means that in each communication round, two fewer clients need to compute their updates, leading to a more efficient path to achieving high accuracy. Note that the images within each superclass were re-clustered based on visual similarity using t-distributed stochastic neighbor embedding, leading to a non-IID distribution.
@@
Figure 7: The plot shows the moving average of test accuracy on ImageNet (with a window size of 5) across 100 communication rounds for three client selection methods: Random selection, training loss-based selection, cluster selection with agglomerative clustering, each with their corresponding parameters.
@@
Finally, we measured the time required to select clients for participation on a local computer, specifically when choosing 10 out of 16 clients for training. The results show that agglomerative clustering is only slightly slower than random selection, with a selection time of 0.0049 seconds compared to 0.0027 seconds for random selection. However, HDBSCAN requires significantly more time, taking 0.0444 seconds. Despite this, the communication time for transmitting weights in a network will far exceed these client selection times, especially in a cross-silo setup with a relatively low number of clients making the additional computation time for clustering methods less impactful in the overall process.
@@ IV Conclusion and Outlook
This paper explored the complexities and strategies associated with client selection and integrating new clients into federated learning systems. We highlighted how FL stands as a paradigm shift from traditional centralized machine learning models, emphasizing its potential to leverage data privacy and enhance model training without compromising data security. Sophisticated client selection strategies can optimize the learning process and ensure that the models developed are more comprehensive and representative of diverse data scenarios, enhancing their real-world applicability.
The client selection strategies discussed in this paper provide a basis for practical applications in FL, particularly in sectors like manufacturing where data sensitivity is paramount. By continuing to refine these methods, we can improve the efficiency and effectiveness of FL systems, ensuring they remain resilient and adaptable in dynamic environments with continuously evolving data and client landscapes. As we move forward, it will be crucial to focus on developing more nuanced client selection and clustering techniques that can further enhance the performance and fairness of FL systems and extend it to other architectures. Our future work will explore an adaptive adjustment of the number of clusters to respond to changing conditions and the inclusion of new clients.

### v2

# Accepted to be published in: The 2nd IEEE International Conference on Federated Learning Technologies and Applications (FLTA24)

@@ II State of the Art
Figure 1: Illustration of the federated learning process [19].
@@ III Methodology
First, we examined how clustering approaches, which were primarily demonstrated on MNIST and minimal examples, can be transferred to real-world datasets and state-of-the-art neural networks, such as EfficientNet or DenseNet. Subsequently, we evaluated various clustering methods and selected the most effective one for further analysis. Initial testing was conducted on a somewhat simpler, proprietary dataset (see section III-A) to establish a baseline understanding, which was then validated on a subset of the more complex ImageNet dataset. This approach allowed us to assess the generalizability and robustness of the selected clustering method across different types of data and neural network architectures.
Even relatively small neural networks typically contain millions of trainable parameters. When fine-tuning on custom datasets, it is common practice to adapt only the final layers to the specific data and the required number of output classes. Consequently, analyzing only the final layer’s activations is often sufficient for effective clustering, as it captures the most relevant features learned during fine-tuning. In this case, with a shape of (n, 1280) corresponding to n classes in EfficientNet, this approach results in a small sample size with high dimensionality, which presents unique challenges for clustering.
@@ III-A Dataset
Our Dataset consists of cabins of miniature trucks, where the parts can be 3D printed, milled, or bought separately. See Figure 2 for two exemplary images from the quality inspection module. As the product is assembled on demand in lot size one, many combinations are possible. Previous publications proved that one advantage of FL is the higher generalizability [43], therefore allowing us to take many pictures away from the production line, where a higher variability of image date is easier to obtain and the risk of interfering with the production process is avoided. For testing purposes, we split the data set so that each client has only one color and one type of windshield, in addition to the class without windshield for every client. The distribution can be obtained from Table I: Classes 1 to 4 are evenly distributed at approximately 12%, whereas class 0 is the predominant category, comprising nearly 50% of the total distribution. In contrast, class 5 is underrepresented, accounting for only about 5%. For some tests, not all clients take part in the federation from the beginning and simulate newcomers that join with new data.
@@ III-B Usability Study on Clustering
A pretrained EfficientNet B4 [44] was used as the frozen backbone, with one fully connected layer replaced to match the number of classes. The training was done with SGD as the optimizer with a learning rate of 0.001, 200 communication rounds and 5 local epochs. Using Pytorch, the preliminary examinations for clustering were carried out with our implementation to exclude any other influences (e.g. calculating the global model manually based on client’s weight files).
To determine which clustering method is most appropriate, three methods that are well-suited for a large number of data points and clusters were selected: Agglomerative clustering [45], HDBSCAN (Hierarchical Density-Based Spatial Clustering of Applications with Noise) [46] and BIRCH (Balanced Iterative Reducing and Clustering using Hierarchies) [47]. Figure 3 presents sample results from a larger set of tests that were conducted. The optimal parameters, such as the distance metric, for each method were determined through grid search on the weights from previous runs in which all clients participated. We observe that with agglomerative clustering, accuracy increases rapidly, particularly during the first 10 rounds, and then remains stable while continuing to improve gradually. All other methods exhibit significantly higher fluctuations and, except for BIRCH (n=5), even achieve lower accuracy than agglomerative clustering. It should be noted that BIRCH occasionally even clearly surpasses it. The high fluctuation in BIRCH’s performance could be due to its sensitivity to parameters and the high dimensionality combined with a small sample size. BIRCH is generally designed for larger datasets, and when applied to a small dataset with high dimensional features, it may struggle to form stable clusters, leading to the observed variability in test accuracy. HDBSCAN consistently performed worse in comparison.
@@ III-C Evaluation of Client Selection Strategies
In order to compare different strategies, the clients as listed in Table II were selected for training and the data of the excluded clients were used for testing, whereby random images were selected to create a similar distribution of classes, e.g. to avoid over-representing class 0. The total number of clients was 22, 16 clients correspond to the setup shown in Table II. In order to simulate the complete absence of a class, the last client with class 5 was omitted from the training in some tests, resulting in 15 clients. The configuration for each run was: EfficientNet with SGD as an optimizer with learning rate 0.001, 200 communication rounds with 5 local epochs each. As a baseline, random sampling of approximately twenty and fifty percent (rounding up to the next integer number) of all participating clients was done. Between a sampling rate of 50% and 20% of all 22 clients, there is no noticeable difference, supporting the findings of [18] that saturation occurs at a certain point, and additional client participation no longer results in improvement.
@@
Figure 5 illustrates the test accuracy over 200 communication rounds for four client selection methods: Random selection with 5 clients participating each round, loss-based selection of the 5 clients with the highest training loss, and agglomerative clustering with k=3 and k=5, corresponding to the number of clusters to be identified. The moving average with windows size 5 is used for clearer visualization. Agglomerative clustering with k=5 consistently outperforms the other methods, showing the highest and most stable accuracy throughout. Initially, it demonstrates a rapid increase in accuracy, maintaining superior performance across all phases. Agglomerative clustering with k=3 and random selection yield similar results, with k=3 exhibiting more fluctuations but eventually aligning with random selection. The training loss method, despite early instability, eventually achieves comparable accuracy, though with more fluctuations. Overall, clustering with k=5 offers the greatest accuracy and stability, while non-randomized methods provide better and more controlled integration of new clients.
@@
Figure 4: Comparison of client participation frequencies across different selection methods: The chart displays the number of participations for each client using three selection strategies: Random selection, highest training loss and cluster selection with k=3.
@@
Figure 5: The plot shows the moving average of test accuracy across 200 communication rounds for three client selection methods: Random selection, training loss-based selection, and agglomerative clustering, each with their corresponding parameters.
To ensure that cluster-based client selection is effective beyond our dataset, we conducted further experiments using a subset of ImageNet. This subset includes three superclasses: animals, tools, and vehicles, each containing multiple subclasses (for instance, various classes for cats, dogs, and birds under the animals’ superclass). Throughout all runs, the mapping of clients to datasets remained consistent, as did the randomization seed, such as when adding noise to the initially loaded ImageNet weights. The results, shown in Figure 6, demonstrate that clustering strategies with any amount of clusters outperform the random and training loss-based strategies, particularly during the first 25 communication rounds. This analysis was conducted using the same hyperparameters as before, highlighting the effectiveness of clustering approaches in improving model accuracy, especially in the early stages of training, but also proofing that ImageNet is more complex and needs more fine-tuning to achieve a higher accuracy. Note, that clustering with k=3 outperforms random selection with n=5, meaning that in each communication round, two fewer clients are required to compute their updates, leading to a more efficient path to achieving high accuracy. It is important to highlight that the images within each superclass were re-clustered based on visual similarity using t-distributed stochastic neighbor embedding, leading to a non-IID distribution.
Figure 6: The plot shows the moving average of test accuracy on ImageNet across 100 communication rounds for three client selection methods: Random selection, training loss-based selection, and agglomerative clustering, each with their corresponding parameters.
III-D Computational Overhead
@@
Finally, we measured the client selection time on a local computer, as consistent hardware configurations cannot be guaranteed throughout all tests when using a high-performance cluster. When selecting 10 out of 16 clients for training—already a higher ratio compared to our other setups—the results indicate that agglomerative clustering is only slightly slower than random selection, with selection times of 0.0049 seconds and 0.0027 seconds, respectively. However, HDBSCAN requires significantly more time, taking 0.0444 seconds. Despite this, the communication time for transmitting weights in a network will far exceed these client selection times, especially in a cross-silo setup with a relatively low number of clients making the additional computation time for clustering methods less impactful in the overall process.
@@ IV Conclusion and Outlook
In this study, we examined various client selection methods within the context of federated learning to address the challenges of integrating new clients with diverse and potentially non-IID data distributions. Our analysis demonstrated that clustering techniques, particularly agglomerative clustering, consistently outperformed random and loss-based client selection methods, offering both higher and more stable accuracy across communication rounds. This improvement is especially pronounced in the early stages of training, where efficient client selection plays a crucial role in accelerating the learning process while maintaining high model performance. The success of clustering-based methods underscores their potential in managing the complexities of data heterogeneity and enhancing the overall efficiency of FL systems.
Future research will focus on further refining clustering-based client selection strategies to better accommodate the dynamic nature of real-world FL environments, where clients may join or leave the system unpredictably. Additionally, the consideration of complex datasets like ImageNet, which has only been touched on in this study, is explored in greater depth in ongoing research and will be crucial to ensure their applicability across various domains. Another promising direction is the development of hybrid approaches that combine the strengths of different selection methods to optimize both accuracy and computational efficiency. As FL continues to evolve, addressing these challenges will be key to its successful adoption in production environments, particularly in industries where data privacy and diversity are paramount.